OpenStudy (anonymous):

A drawer contains 3 tank tops, 6 short-sleeve shirts, 4 three-quarter sleeve shirts, and 5 long-sleeve shirts. What is the probability of randomly choosing a tank top and then a three-quarter sleeve length, if you replace the tank top?

OpenStudy (anonymous):

@heretohelpalways

OpenStudy (anonymous):

I am here, again

OpenStudy (anonymous):

lol hiiii:)

OpenStudy (anonymous):

add the numbers up

OpenStudy (anonymous):

=18

OpenStudy (anonymous):

Is this in fraction form?

OpenStudy (anonymous):

so 3/18*4/18

OpenStudy (anonymous):

12/18

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

1/27

OpenStudy (anonymous):

what are your options?

OpenStudy (anonymous):

1/27 is an option:D

OpenStudy (mathstudent55):

\(\dfrac{3}{18} \times \dfrac{4}{18} = \dfrac{3 \times 4}{18 \times 18} = \dfrac{ \cancel{3} \times \cancel{4}2}{\cancel{18}6 \times \cancel{18}9} = \dfrac{2}{54} = \dfrac{1}{27}\)